# The Relative Entropy of Expectation and Price

### v4

@@ III.5 Heston model
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 1,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 14: In these examples of data-driven models, the underlying payoff $Q:=q+dq$ is modelled in the expectation measure as a uniform variable on 10,000 samples from a standard normal variable. In each row, the price measure is supported on the same samples, adjusting the probabilities to match the target underlying price $q$ with minimum entropy relative to the expectation measure. No adjustment is necessary when $q=0$, and the probabilities are downward/upward sloping when $q$ is negative/positive, biasing the distribution in the direction of the target underlying price.
      The graphs plot the delta and price per unit notional as a function of the strike $k$ of a digital option with unit payoff $P:=1_{Q>k}$ for a range of notionals $\lambda$, highlighting the central plot for the mid-pricing case $\lambda=0$. Risk aversion is fixed at one, $\alpha=1$, and the implied funding rate is set to zero, $r=0$.
III.6 Data-driven model
In this section, assume that the model is supported on a finite set of states with uniform probability in the expectation measure.
Assumption III.6 (Data-driven model).
The model is supported on a finite set of states that are equally likely in the expectation measure. In the $n$th state, the underlying return is $dq_{n}$ and the final derivative price is $P_{n}$.
For this data-driven approach, the expectation measure is represented as a uniform distribution on a finite set of states, which could for example be sampled from a generative model trained on historical market data. The calibration, hedge and price conditions are discretised as:
00 0 $=\sum\nolimits_{n}\gamma_{n}(dq_{n}-qr\,dt)$ (74)
00 0 $=\sum\nolimits_{n}\gamma_{n}(dq_{n}-q(r+\alpha s)\,dt)\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$
$p$ $=-\frac{1}{\alpha}\log\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$
where the price measure is supported on the same set of states but with non-uniform probabilities:
$\gamma_{n}:=\frac{\displaystyle\exp[-\phi\cdot dq_{n}]}{\displaystyle\sum\nolimits_{m}\exp[-\phi\cdot dq_{m}]}$ (75)
These conditions are accurately and efficiently solved using the Newton-Raphson scheme.
As demonstrated in the examples of figures 9, 10, 11, 12, 13 and 14, the hedge and price conditions can be solved exactly on the assumption of the discrete expectation measure, but the solution is impacted by the choice. In the example with 100 sample points, the support of the measure does not extend beyond the interval (-2.5,2.5), and there can be no option value outside this range of strikes. Increasing the number of points brings the solutions closer to the limit of the standard normal variable used to generate the samples. The support of the expectation measure plays a crucial role in determining hedge and price and also the extent to which unhedged risks impact the spread of prices. Unrealistic assumptions on the support mask these model risks and can hinder their quantification.
There are situations where it is useful to reframe the calibration, hedge and price conditions as optimisation problems. The calibration condition is solved for the unit optimal portfolio $\phi$ via the optimisation problem:
$\phi=\operatorname*{arg\,min}\nolimits_{\psi}\!\left\{\sum\nolimits_{n}\exp[-\psi\cdot dq_{n}]:\psi\cdot q=0\right\}$ (76)
The hedge and price conditions are then solved simultaneously for the hedge portfolio $\delta$ and the initial derivative price $p$ via the optimisation problem:
@@ III.6 Data-driven model
$\delta$ $=\operatorname*{arg\,min}\nolimits_{\epsilon}\!\left\{\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\epsilon\cdot dq_{n})]:\epsilon\cdot q=p\right\}$
$p$ $=-\frac{1}{\alpha}\log\sum\nolimits_{n}\gamma_{n}\exp[-\alpha(P_{n}-\delta\cdot dq_{n})]$ (77)
These expressions are amenable to machine learning methods and are the starting point for a variety of applications in deep hedging. Expressed as optimisation problems, additional features such as frictional trading costs can be included in the program, and large datasets can be accommodated by using universal approximations for the hedging strategy.
The examples of figures 9, 10, 11, 12, 13 and 14 demonstrate the reach of the entropic framework. With no theory to support development beyond the information that the data expresses, data-driven models need a more general foundation that does not depend on unrealistic assumptions such as complete markets, continuous hedging or risk-free funding. Advocates of deep hedging address this by replacing dynamic risk neutralisation with dynamic risk optimisation, using convex risk metrics to assess strategic performance. While they suffer from lack of explainability, data-driven models can uncover complex relationships between economic variables that may be hard to capture theoretically.
IV Market and model risk
Returning to the general pricing framework, with the expectation measure $\mathbb{E}$ the calibration conditions:
00 0 $=\phi\cdot q$ (78)
00 0 $=\mathbb{E}^{\phi}[dq]-qr\,dt$
$p$ $=\delta\cdot q$ (79)
$p$ $=\mathbb{E}^{\phi}[\alpha][P-\delta\cdot dq]$
are solved to generate the implied funding spread $s$, the hedge portfolio $\delta$ and the initial price $p$ for the derivative security. While the hedge neutralises the underlying component of the derivative risk, when the market is incomplete there are unhedged risks that are sensitive to the choice of expectation measure. The entropic pricing framework thus distinguishes between the market and model risks in the derivative security.
$\partial\phi$ $=\mathbb{V}^{\phi}[dq]^{-1}(\mathbb{V}^{\phi}[dq,W]-q\bar{r}\,dt)+O[\epsilon]$
$\bar{\mathbb{E}}^{\bar{\phi}}[X]-\mathbb{E}^{\phi}[X]=$ (82)
$\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][X]-\mathbb{E}^{\phi}[\alpha][X]=$
for the undetermined variable $X$.
$\bar{p}$ $=\bar{\delta}\cdot q$ (83)
$\bar{p}$ $=\bar{\mathbb{E}}^{\bar{\phi}}[\alpha][P-\bar{\delta}\cdot dq]$
$\partial p=\frac{1}{1+(r+\alpha s)\,dt}\times$ (84)
V Entropic XVA
@@
Refer to caption Refer to caption
$\lambda=0$ and $c=\infty$ $\lambda=0$ and $c=0.5$
Refer to caption Refer to caption
$\lambda=1$ and $c=\infty$ $\lambda=1$ and $c=0.5$
Refer to caption Refer to caption
$\lambda=2$ and $c=\infty$ $\lambda=2$ and $c=0.5$
Figure 18: In this example, two counterparties have differing expectations for the future state. The first counterparty models the state as a uniform variable with zero mean and unit standard deviation. The second counterparty models the state as a normal variable with mean $\mu$ and unit standard deviation. The trade is then a call option on the state with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
@@
V.1 Margined derivatives
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Revisiting the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a derivative trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The derivative is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in figures 17, 17 and 17, net of all considerations the capital returned from the derivative is:
@@ V.2 Entropic XVA
where $P:=p+dp$ is the final derivative price and the margins are $m_{1}:=c_{1}-p$ and $m_{2}:=c_{2}+p$ for the two counterparties. The first term is the discounted entropy-adjusted mean in the price measure of the final derivative price, capped and floored by the respective margins to generate the credit and debit valuation adjustments. The second term then adjusts the initial price by the differential between the returns on margin from the underlying security versus the margin account, creating the funding valuation adjustment.
V.3 Entropic margin
@@
$m_{1}$ $:=c_{1}-p\geq-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (90)
$m_{2}$ $:=c_{2}+p\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]$
where $P:=p+dp$ is the contractual settlement of the derivative security. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures.
@@ VI Literature review
Closer consideration of market incompleteness led to the development of alternative pricing principles based on the dynamic optimisation of convex risk metrics, as presented in [8, 9, 10]. With foundations that extend well beyond the domain of applicability of risk-neutral pricing, these principles enabled the development of data-driven methods including the pioneering work on deep hedging by Buehler and others [11, 12]. As a recent practical example, Oya [13] applies these methods to explain the mark-to-market valuation adjustment for Bermudan swaptions.
Entropy methods have found application across the range of mathematical finance, including portfolio optimisation [14] and derivative pricing [15, 16] – see also the review essay [17] which includes further references. The article [18] by Frittelli proposes the minimal relative entropy measure as a solution to the problem of pricing in incomplete markets, and links this solution with the maximisation of expected exponential utility. The entropic solution is connected with risk optimisation in the introduction [19], which this essay extends with the adoption of the entropic risk metric. As a measure of disorder, entropy performs a role similar to variance but is better suited to the real distributions of market returns. Proponents of the use of entropy justify the approach by appeal to the modelling of information flows in dynamical systems and the analogy with thermodynamics.
@@
It is remarkable that the principle of dynamic convex risk optimisation enables the consistent and simultaneous application of models across electronic trading algorithms, for pricing and hedging exotic derivatives, and to determine economic and regulatory capital. In this context, the entropic risk metric is promoted for its simple implementation as an extension of the risk-neutral framework, which supports integration into legacy platforms.

### v5

@@ V Discrete models
Call Option (100 states) Call Option (1,000 states) Call Option (10,000 states)
Figure 12: In these studies of discrete models on quantum information, the final hedging price is uniformly distributed in the expectation measure on $n=2$ values $-1$ and $1$, and is decorrelated from the final funding price by applying the rotation $\theta=0$, $\pi/5$ and $\pi/4$ to its eigenbasis. Entropic risk optimisation identifies the mid and bid-offer for the price and the hedge ratio for a range of notionals $\lambda$.
Entropy is a universal concept in information theory that is ideally suited to a wide range of applications. In this section, the entropic risk metric is applied to risk optimisation in discrete models of classical and quantum information. The core premise of the discrete model is that the indeterminate variable has a spectrum of possible eigenvalues, which may include duplicates, that are uniformly distributed in the expectation measure. Classical or quantum models are then characterised by the existence or absence of a common basis of eigenstates supporting the eigenvalues.
The indeterminate variable $X$ is associated with its spectrum $(X_{1},\ldots,X_{n})$ and the expectation measure assumes these outcomes are equally likely. Represent the variable as a diagonal matrix with its eigenvalues arrayed on the diagonal. The mean of $X$ is the average eigenvalue, expressed as the normalised trace of its matrix:
$\mathbb{E}[X]=\frac{1}{n}\operatorname{tr}[X]$ (87)
Significantly, this definition extends to arbitrary self-adjoint matrices, where the mean, defined to be the average eigenvalue, is equivalently evaluated as the normalised trace. Classical information is characterised by the commutativity of its indeterminate variables; without loss of generality, the variables are simultaneously diagonalised and the expectation measure is uniformly distributed on the common set of diagonalising eigenstates. This restriction is lifted for quantum information, where non-commutativity is embraced as an additional source of uncertainty with remarkable consequences for the quality of model fitting.
Functional calculus extends to self-adjoint matrices by applying the transformation directly to the eigenvalues in the diagonalised representation. The entropy-adjusted mean of $X$ is then defined:
$\mathbb{E}[\alpha][X]=-\frac{1}{\alpha}\log\!\left[\frac{1}{n}\operatorname{tr}[\exp[-\alpha X]]\right]$ (88)
When optimising the strategy with this target function, care must be taken to manage the potential complications of non-commutativity. For example, the perturbation $\partial(X^{2})=(\partial X)X+X(\partial X)$ cannot be simplified further when $X$ and $\partial X$ do not commute. Fortunately, the trace of this expression can be simplified, thanks to cyclicity of the trace, and in general the perturbation has the handy simplification:
$\partial\operatorname{tr}[f[X]]=\operatorname{tr}[f^{\prime}[X]\,\partial X]$ (89)
The calibration, hedge and price conditions are then derived from entropic risk optimisation as:
00 0 $=\operatorname{tr}[(dq-qr\,dt)\exp[-\alpha\omega\cdot dq]]$ (90)
00 0 $=\operatorname{tr}[(dq-q(r+\alpha s)\,dt)\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]$
@@
$p$ $=-\frac{1}{\alpha}\log\!\left[\frac{\operatorname{tr}[\exp[-\alpha(P+(\omega-\delta)\cdot dq)]]}{\operatorname{tr}[\exp[-\alpha\omega\cdot dq]]}\right]$
These conditions are explored in the studies of figures 9, 10, 11 and 12, assuming an underlying market that contains a funding security with initial price $b=1$ and final price $B=b+db$ satisfying $\operatorname{tr}[db]=0$ and a hedging security with initial price $q$ and final price $Q=q+dq$ satisfying $\operatorname{tr}[Q]=0$. The model is developed for the final derivative price expressed as a function $P=f[Q-kB]$ of the spread $Q-kB$ at strike $k$.
V.1 Classical study
The classical studies of figures 9, 10 and 11 assume $B=1$ and $Q$ is a diagonal matrix with eigenvalues $Q_{i}$ sampled from a standard normal variable, centralised to have zero sample mean. The matrix $P=f[Q-kB]$ is then diagonal with eigenvalues $P_{i}:=f[Q_{i}-k]$.
The calibration condition:
$0=\sum_{i=1}^{n}(Q_{i}-q)\exp[-\alpha\omega Q_{i}]$ (91)
is solved for the optimal portfolio $\omega$, which generates the eigenstate probabilities:
$\gamma_{i}:=\frac{\displaystyle\exp[-\alpha\omega Q_{i}]}{\displaystyle\sum\nolimits_{j=1}^{n}\exp[-\alpha\omega Q_{j}]}$ (92)
in the price measure. The hedge condition:
$0=\sum_{i=1}^{n}\gamma_{i}(Q_{i}-q)\exp[-\alpha(P_{i}-\delta Q_{i})]$ (93)
is then solved for the hedge portfolio $\delta$. Using these solutions, the price condition:
$p=\delta q-\frac{1}{\alpha}\log\!\left[\sum_{i=1}^{n}\gamma_{i}\exp[-\alpha(P_{i}-\delta Q_{i})]\right]$ (94)
generates the initial derivative price $p$.
Three cases are considered in the studies with $n=$100$$ (left column), $1000$ (middle column) and $10000$ (right column), for the call option $f[x]=x^{+}$ (top two rows) and the digital option $f[x]=1_{x>0}$ (bottom two rows). The derivative security cannot be exactly replicated by the funding and hedging securities, and entropic risk optimisation compensates for this model risk with a bid-offer around the mid price and a skew to the hedge ratio. In the study with $n=$100$$ samples, the support is contained within the interval (-2.5,2.5), and there can be no option value outside this range of strikes. Increasing the number of samples widens the support and brings the solution closer to the limit of the standard normal variable.
The support of the expectation measure thus plays a crucial role in determining hedge and price and also the extent to which unhedged risks impact the bid-offer. Entropic pricing does not allocate probability beyond the support of the expectation measure, and unrealistic assumptions on the support can mask model risks.
V.2 Quantum study
The quantum studies of figure 12 assume $q=0$ and $B$ and $Q$ are matrices with $n=2$ eigenvalues:
$B$ $=1+\epsilon\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$ (95)
$Q$ $=\begin{bmatrix}\cos[\theta]&-\sin[\theta]\\ \sin[\theta]&\cos[\theta]\end{bmatrix}\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}\begin{bmatrix}\cos[\theta]&\sin[\theta]\\ -\sin[\theta]&\cos[\theta]\end{bmatrix}$
The gap $\epsilon$ separates the eigenvalues of the funding security and the angle $\theta$ rotates the eigenbasis of the hedging security. The matrix $P=f[Q-kB]$ is obtained by first diagonalising the matrix $Q-kB$ and then applying the function $f$ to its eigenvalues:
$P$ $=$ (96)
$\begin{bmatrix}\cos[\psi]&-\sin[\psi]\\ \sin[\psi]&\cos[\psi]\end{bmatrix}\begin{bmatrix}P_{+}&0\\ 0&P_{-}\end{bmatrix}\begin{bmatrix}\cos[\psi]&\sin[\psi]\\ -\sin[\psi]&\cos[\psi]\end{bmatrix}$
with:
$\psi$ $:=\frac{1}{2}\arctan\!\left[\frac{\sin[2\theta]}{\cos[2\theta]-\epsilon k}\right]$ (97)
$Q_{\pm}$ :=±$:=\pm\frac{\sin[2\theta]}{\sin[2\psi]}$
$P_{\pm}$ $:=f[Q_{\pm}-k]$
The non-linear relationship between a matrix and its eigenvalues is exploited to create a complex dependence of the option payoff on the strike. Fine tuning the rotation of the eigenbases can then be used to fit the implied volatility smile.
The calibration condition is trivially solved with $\omega=0$, and the hedge and price conditions become:
$\delta$ $=\hat{P}\cos[2(\psi-\theta)]-\epsilon p\cos[2\theta]$ (98)
$p$ $=\bar{P}-\frac{1}{\alpha}\log\cosh[\alpha(\hat{P}\sin[2(\psi-\theta)]+\epsilon p\sin[2\theta])]$
where $\bar{P}:=(P_{+}+P_{-})/2$ and $\hat{P}:=(P_{+}-P_{-})/2$. The price condition is first solved for the initial derivative price $p$. Using this solution, the hedge condition generates the hedge portfolio $\delta$.
Three cases are considered in the studies with $\theta=0$ (left column), $\pi/5$ (middle column) and $\pi/4$ (right column), with $\epsilon=1-\cos[2\theta]$ in each case, for the call option $f[x]=x^{+}$. Contrasting with the classical studies where thousands of eigenstates are needed for acceptable convergence, the inclusion of non-commutativity between the final funding and hedging prices generates a credible model for the initial derivative price with as few as two eigenstates.
With two securities available for funding and hedging, the derivative security is exactly replicated in the classical binomial model with zero rotation. Heisenberg uncertainty breaks market completeness, and entropic risk optimisation creates a compensating bid-offer around the mid price in the quantum binomial model when the rotation is non-zero. At the maximum rotation $\theta=\pi/4$, the implied volatility smile has continuous support on the full line, a remarkable property for a model whose final underlying prices are binomial.
VI Entropic XVA
@@
Refer to caption Refer to caption Refer to caption
$\lambda=0$ and $c=\infty$ $\lambda=1$ and $c=\infty$ $\lambda=2$ and $c=\infty$
Refer to caption Refer to caption Refer to caption
$\lambda=0$ and $c=0.5$ $\lambda=1$ and $c=0.5$ $\lambda=2$ and $c=0.5$
Figure 16: In this study, two counterparties have differing expectations for the future state. The first counterparty models the final underlying price as a uniform variable with zero mean and unit standard deviation. The second counterparty models the final underlying price as a normal variable with mean $\mu$ and unit standard deviation. The trade is then a call option on the final underlying price with zero strike. In each graph, the entropy-adjusted mean return on capital, computed with risk aversion $\alpha=1$, is plotted for both counterparties as a function of the option price per unit notional, including a range of values for the mean $\mu$ assumed by the second counterparty. The first counterparty is fully capitalised against default, and each row includes cases $c=\infty,0.5$ for the capital of the second counterparty. Each column includes cases $\lambda=0,1,2$ for the notional, where the case $\lambda=0$ corresponds to mid-pricing.
@@
VI.1 Margined derivatives
If the true costs of risk management are fully accounted in the return, reckless investments are naturally penalised by their associated margin costs. Returning to the bilateral model outlined in the introduction, suppose that two counterparties deploy capital $c_{1}$ and $c_{2}$ to a derivative trade with initial price $p$. Excess capital is deposited in the margin account with unit price $b>0$. The derivative is settled from the margin account with no additional funds injected, creating the risk that insufficient capital results in default by one of the counterparties.
As shown in figures 15, 15 and 15, net of all considerations the capital returned from the derivative is:
@@ VI.2 Entropic XVA
where $P=p+dp$ is the final derivative price and the margins are $m_{1}:=c_{1}-p$ and $m_{2}:=c_{2}+p$ for the two counterparties. The first term is the discounted entropy-adjusted mean in the price measure of the final derivative price, capped and floored by the respective margins to generate the credit and debit valuation adjustments. The second term then adjusts the initial price by the funding valuation adjustment.
VI.3 Entropic margin
@@
$m_{1}$ $\geq-\operatorname*{ess\,inf}\nolimits_{1}\!\left[\frac{P}{1+db/b}\right]$ (104)
$m_{2}$ $\geq\operatorname*{ess\,sup}\nolimits_{2}\!\left[\frac{P}{1+db/b}\right]$
for the contractual settlement $P$ of the derivative security. Setting these inequalities to equalities generates the minimum capital required to fully guarantee settlement, under the assumptions of the respective expectation measures.
@@ VII Literature review
Closer consideration of market incompleteness led to the development of alternative pricing principles based on the dynamic optimisation of convex risk metrics, as presented in [13, 14, 15]. With foundations that extend well beyond the domain of applicability of risk-neutral pricing, these principles enabled the development of data-driven methods including the pioneering work on deep hedging by Buehler and others [16, 17]. As a recent practical example, Oya [18] applies these methods to explain the mark-to-market valuation adjustment for Bermudan swaptions. Applications of quantum information in option pricing are investigated further in [19], and the technical requirements to extend this approach to continuous-time models are developed in [20].
Entropy methods have found application across the range of mathematical finance, including portfolio optimisation [21] and derivative pricing [22, 23] – see also the review essay [24] which includes further references. The article [25] by Frittelli proposes the minimal relative entropy measure as a solution to the problem of pricing in incomplete markets, and links this solution with the maximisation of expected exponential utility. The entropic solution is connected with risk optimisation in the introductory essay [26], which this essay extends with the adoption of the entropic risk metric. As a measure of disorder, entropy performs a role similar to variance but is better suited to the real distributions of market returns. Proponents of the use of entropy justify the approach by appeal to the modelling of information flows in dynamical systems and the analogy with thermodynamics.
@@
It is remarkable that the principle of dynamic convex risk optimisation enables the consistent and simultaneous application of models across electronic trading algorithms, for pricing and hedging exotic derivatives, and to determine economic and regulatory capital. In this context, the entropic risk metric is promoted for its applicability within a diverse range of information models and its simple implementation as an extension of the risk-neutral framework, which supports integration into legacy platforms.